# Richer Output for Richer Countries: Uncovering Geographical Disparities in Generated Stories and Travel Recommendations

### v1

@@ Abstract
While a large body of work inspects language models for biases concerning gender, race, occupation and religion, biases of geographical nature are relatively less explored. Some recent studies benchmark the degree to which large language models encode geospatial knowledge. However, the impact of the encoded geographical knowledge (or lack thereof) on real-world applications has not been documented. In this work, we examine large language models for two common scenarios that require geographical knowledge: (a) travel recommendations and (b) geo-anchored story generation. Specifically, we study four popular language models, and across about K travel requests, and K story generations, we observe that travel recommendations corresponding to poorer countries are less unique with fewer location references, and stories from these regions more often convey emotions of hardship and sadness compared to those from wealthier nations. 111Upon publication, we will publicly release the data and code required to reproduce our evaluation.
@@ 1 Introduction
Through our analysis, we uncover several geographical disparities, finding that outputs for wealthier countries are more unique and include more geographical entities (Figure 1). For Sub-Saharan African countries, we notice considerably less unique outputs compared to the North American region, with the average difference being about % across all models. Further, stories about poorer countries express considerably more hardship and sadness, with % more narratives depicting hardship for low-income countries compared to high-income countries.
@@ 3 Approach
| Scenario | Attribute | GPT4 | Mixtral 8x7B | LLaMa3 8B | LLaMa3 70B | ||||
| Travel | GDP | Freq | GDP | Freq | GDP | Freq | GDP | Freq | |
| Rec. | Uniqueness | ||||||||
| # Geo-entities | |||||||||
| Absence of Info | — | — | — | — | |||||
| Story | Uniqueness | ||||||||
| Gen. | # Geo-entities | ||||||||
| Emotion (Hardship) | |||||||||
| Emotion (Sadness) | |||||||||
@@ 3.1 Experimental Setup
To capture the global variation in LLM performance, it is crucial to incorporate a diverse array of geographical regions, covering cities, towns, and villages, in our analysis. For this purpose, we use Geonames333https://www.geonames.org/, a community-driven geographical database, comprising an extensive list of location names. We capture all global locations with inhabitants exceeding , to achieve a finer granularity and a more nuanced understanding of global disparities. This surpasses most prior work that largely study geographical biases with only country-level information (§2). At the time of our study, GeoNames contained about K locations. In our experiment, we randomly sample, with replacement, up to locations per country from this larger population totalling about 4,000 locations for 190 countries. For every attribute we quantify, we report averages across random samples (of this data) with different seeds.
@@
We perform our evaluation on four widely used and capable models: GPT- (Achiam et al., 2023), Mixtral B (Jiang et al., 2024), LLaMa B, and LLaMa B (Dubey et al., 2024), to help us cover a broad range of characteristics, including size, training methods, and availability. We set the temperature to to achieve a balance between creativity and coherence.
@@ 3.2 Evaluation
Every location presents a blend of historical, cultural, geographical, and environmental factors, contributing to its unique identity. To reflect this, models need to be aware of these distinctive aspects for different regions. We use the uniqueness measure to capture how distinct the subjective responses for a given location are compared to others. We make a slight modification to the TF-IDF metric to calculate the average rarity of words generated in the response of a location compared to other locations. Terms which score high as per our metric, anecdotally, reflect regional artifacts (e.g., “dosa” for India). We provide further details of the uniqueness metric in Appendix A.2.
@@ 4 Results
We aggregate the (estimated) city-level attributes for every country to observe trends across different countries. Interestingly, the uniqueness scores for India, Italy, Japan, and the United States consistently rank among the highest across all models in both applications. Comparing regions, we observe that travel recommendations for the Sub-Saharan African region are considerably less unique than those for the North American region, with alarming differences of % for GPT-4, % for LLaMa B, % for LLaMa B, and % for Mixtral.
A similar trend is observed in the informativeness of generated stories, wherein most models generate a large number of geographical entities for the United States and the United Kingdom. Problematically, the responses for the North American region include at least double the number of location mentions in stories compared to those from the Sub-Saharan African region across all models.
@@
We examine two models from the LLaMa family, only differing in size, and inquire whether larger models hold more information about global locations, thus potentially reducing disparities. However, our results suggest otherwise as the larger model results in similar correlations.
@@
5 Conclusion
We evaluate language models on two real-world applications of generating travel recommendations and geo-anchored stories, through a geographical lens. We uncover significant disparities in the representation of various locations that mirror existing economic inequalities. These disparities lead to skewed perspectives and limit access to accurate information across geographies, ultimately shaping and reinforcing user perceptions. This study underscores the importance of developing geographically diverse datasets for AI training, to create more equitable and representative models that better serve a wide range of global needs.
6 Limitations
@@
Third, this study primarily focuses on attributes amenable to easy quantification, however, a qualitative study to evaluate the cultural appropriateness of the responses could yield further insights, and provide a complementary perspective.

### v2

@@ Abstract
While a large body of work inspects language models for biases concerning gender, race, occupation and religion, biases of geographical nature are relatively less explored. Some recent studies benchmark the degree to which large language models encode geospatial knowledge. However, the impact of the encoded geographical knowledge (or lack thereof) on real-world applications has not been documented. In this work, we examine large language models for two common scenarios that require geographical knowledge: (a) travel recommendations and (b) geo-anchored story generation. Specifically, we study five popular language models, and across about K travel requests, and K story generations, we observe that travel recommendations corresponding to poorer countries are less unique with fewer location references, and stories from these regions more often convey emotions of hardship and sadness compared to those from wealthier nations. 111The data and code is available at https://github.com/FLAIR-IISc/richer-countries-have-richer-output.
@@ 1 Introduction
Through our analysis, we uncover several geographical disparities, finding that outputs for wealthier countries are more unique and include more geographical entities (Figure 1). For Sub-Saharan African countries, we notice considerably less unique outputs compared to the North American region, with the average difference being about % across all models. Further, stories about poorer countries express considerably more hardship and sadness, with % fewer narratives depicting hardship for high-income countries compared to low-income countries.
@@ 3 Approach
| Scenario | Attribute | GPT4 | Mistral 7B | Mixtral 8x7B | LLaMa3 8B | LLaMa3 70B | |||||
| Travel | GDP | Freq | GDP | Freq | GDP | Freq | GDP | Freq | GDP | Freq | |
| Rec. | Uniqueness | ||||||||||
| # Geo-entities | |||||||||||
| TTR | |||||||||||
| Absence of Info | — | — | — | — | |||||||
| Story | Uniqueness | ||||||||||
| Gen. | # Geo-entities | ||||||||||
| TTR | |||||||||||
| Hardship | |||||||||||
| Sadness | |||||||||||
@@ 3.1 Experimental Setup
To capture the global variation in LLM performance, it is crucial to incorporate a diverse array of geographical regions, covering cities, towns, and villages, in our analysis. For this purpose, we use Geonames333https://www.geonames.org/, a community-driven geographical database, comprising an extensive list of location names. We capture all global locations with inhabitants exceeding , to achieve a finer granularity and a more nuanced understanding of global disparities. This surpasses most prior work that largely study geographical biases with only country-level information (§2). At the time of our study, GeoNames contained about K locations. To make the analysis computationally feasible, we randomly sample, with replacement, up to locations per country from this larger population totalling about locations for countries. For every attribute we quantify, we report averages across such random samples (with different seeds) to infer patterns across all locations.
@@
We perform our evaluation on five widely used and capable models: GPT- (Achiam et al., 2023), Mistral 7B (Jiang et al., 2023), Mixtral B (Jiang et al., 2024), LLaMa B, and LLaMa B (Dubey et al., 2024), to help us cover a broad range of characteristics, including size, training methods, and availability. We set the temperature to to achieve a balance between creativity and coherence.
@@ 3.2 Evaluation
Every location presents a blend of historical, cultural, geographical, and environmental factors, contributing to its unique identity. To reflect this, models need to be aware of these distinctive aspects for different regions. We use the uniqueness measure to capture how distinct the subjective responses for a given location are compared to others. We make a slight modification to the TF-IDF metric to calculate the average rarity of words generated in the response of a location compared to other locations. Terms which score high as per our metric, anecdotally, reflect regional artifacts (e.g., “dosa” for India). Additionally, we also compute the Type-Token Ratio (TTR), the ratio of unique words to total words, which captures the lexical diversity of these responses. Past efforts have effectively used the TTR metric to measure richness in vocabulary in various contexts (Balestrucci et al., 2024; Morris et al., 2023). We provide further details of the uniqueness metric in Appendix A.2.
@@ 4 Results
We aggregate the (estimated) city-level attributes for every country to observe trends across different countries. Interestingly, the uniqueness scores for India, Italy, Japan, and the United States consistently rank among the highest across all models in both applications. Comparing regions, we observe that travel recommendations for the Sub-Saharan African region are considerably less unique than those for the North American region, with alarming differences of % for GPT-4, % for LLaMa B, % for LLaMa B, % for Mistral B, and % for Mixtral B .
A similar trend is observed in the informativeness of generated stories, wherein most models generate a large number of geographical entities for the United States and the United Kingdom. Problematically, the responses for the North American region include around double the number of location mentions in stories compared to those from the Sub-Saharan African region across all models.
@@
We examine two models each from the LLaMa and Mistral family of models. The key purpose of including models with different sizes is to evaluate whether larger models (which hold more information) exhibit fewer geographical disparities. However, our results suggest otherwise as larger models result in similar correlations.
@@
5 Limitations and Future Work
@@
Third, this study primarily focuses on attributes that are easy to quantify. Future work could consider conducting qualitative studies to evaluate the cultural appropriateness of the responses, which could yield further insights and provide a complementary perspective. Lastly, future work could also expand our evaluation to measure the relevance, relatability and factuality of cultural artifacts within geo-anchored responses, by leveraging existing resources such as CUBE (Kannen et al., 2024) and WorldCuisines (Winata et al., 2024).
6 Conclusion
We evaluate language models on two real-world applications of generating travel recommendations and geo-anchored stories, through a geographical lens. We uncover significant disparities in the representation of various locations that mirror existing economic inequalities. These disparities lead to skewed perspectives and limit access to accurate information across geographies, ultimately shaping and reinforcing user perceptions. This study underscores the importance of developing geographically diverse datasets for model training, to create more equitable and representative models that better serve a wide range of global needs.
Acknowledgements
We would like to thank the anonymous reviewers. Our work benefitted from discussions with Sunayana Sitaram, Mitesh Khapra, Shachi Dave, Partha Talukdar, Mansi Gupta, Preethi Seshadri and Abhipsa Basu. We also appreciate Navreet Kaur for her suggestions on the prompt templates. This work was supported in part by the AI2050 program at Schmidt Sciences (Grant G-24-66186). Additionally, DP is grateful to Adobe Inc., Google Research, and the Kotak IISc AI-ML Centre (KIAC) for generously supporting his group’s research.